OpenStudy (anonymous):

what is the factor by grouping of this trinomial 50v^2+40vt+8t^2

OpenStudy (anonymous):

I am not exactly sure what you asking by factor? But I do know that:\[50v^2+40vt+8t^2=(\sqrt{50}*v+\sqrt8*t)^2=(5\sqrt2*v+2\sqrt2*t)^2\]

OpenStudy (anonymous):

Is that what you mean by grouping?

OpenStudy (anonymous):

yes it is thank you